OpenStudy (anonymous):

Y=9x^2+126x-74. Find min?

OpenStudy (gorv):

dy/dx=18x+126=0 x=-126/18 x=-7 find d^2y/dx^2 =18 second derivative >0 for x=-7 so function will be minimum at x=-7 y min=9*(-7)^2 +126*(-7)-74 =-515 @Mrhnodri

OpenStudy (anonymous):

Thank you :)